OpenStudy (anonymous):

The table below represents a function. x y 0 1 2 5 5 26 7 50 Which equation makes the table true? f(x) = 2x + 1 f(x) = x 2 + 1 f(x) = 2x - 1 f(x) = x 2 - 1

OpenStudy (mrnood):

ok - the only way to do this is by trial and error look at the first equation y= f(x) = 2x+1 Take the pairs from the table and put them in the equation so: 1 = 2*0+ 1 (true) 5= 2*2 + 1 (true) 26 = 2*5 + 1 (FALSE) SO the equation is not true for all the points in th table therefore a is not the answer DO the same for the other 3 equations = one of them will give true answers fro all the pairs in the table..

OpenStudy (anonymous):

ok thnx